OpenStudy (anonymous):

PLEASE HELP:In the below system, solve for y in the first equation. x – 3y = –6 2x – 7y = 10 I got y=-22 but the answers are one thirdx + 2 negative one thirdx + 6 –x + 2 negative one thirdx + 2

OpenStudy (zale101):

your correct :)

OpenStudy (anonymous):

but its not one of the choices :/ thats whhy im confused

OpenStudy (zale101):

what are the choices?

OpenStudy (anonymous):

1/3x+2 -1/3x+6 -x+2 -1/3x+2

OpenStudy (anonymous):

would it be 1/3x+2?

OpenStudy (zale101):

I think so too

OpenStudy (anonymous):

Im going to put that I dont see how it could be anything else

OpenStudy (zale101):

Let's use the first equation and subtract x from both sides

OpenStudy (zale101):

x-3y=6 3y=6−x y=6−x/3

OpenStudy (anonymous):

x-3y=-6 x+6=3y 3y=x+6 \[y=\frac{ 1 }{3}\left( x+6 \right)=\frac{ 1 }{3 }x+2\]